OpenStudy (anonymous):

One model of Earth's population growth is P(t)=64÷(1+11e^(-.08t)), where t is measured in years since 1990, and P is measured in billions of people. Which of the following statements are true? Check all that apply. A. The population of Earth is growing at a rate of just under 8% per year. B. In 1991, there were 5.74 billion people, according to this model. C. The carrying capacity of Earth is 64 billion people. D. The population of Earth will grow exponentially for a while but then start to decrease.

OpenStudy (anonymous):

@campbell_st no idea?

OpenStudy (campbell_st):

well the equation doesn't make a lot of sense to me... if I read it correctly its \[P(t) = \frac{(1 + \frac{11}{e^{0.08t}})}{64}\]

OpenStudy (anonymous):

it looks like this..

OpenStudy (campbell_st):

which seems to indicate exponential decay... rather than growth

OpenStudy (campbell_st):

ok... oops... my mistake... so t = 1

OpenStudy (anonymous):

Im thinking C is a possible answer but I need at least one more answer

OpenStudy (campbell_st):

ok... so looking at it. B is correct... when I substituted t = 1 (representing 1 year since 1990) I got a population P(1) = 5.73771 billion people

OpenStudy (anonymous):

population being 5.7 would make C wrong then?

OpenStudy (campbell_st):

I graphed the curve and D seems to make sense

OpenStudy (anonymous):

so it'll end up being B and D ?

OpenStudy (campbell_st):

the graph flattens out at 64 billion... this is because as t approaches infinity.... e^(-0.08t) approaches zero... so 11*(e^(-0.08t) will approach zero leaving 64/1 = 64... so C seems right...

OpenStudy (anonymous):

hmmm so what would be the final two answers or are you pretty sure B, C and D make sense ?

OpenStudy (campbell_st):

looking at A then t = 0.. the population is 64/12 = 5.333... billion then 1.08 x 5.3333 = 5.76... so 1991 or P(1) is approx 5.76... close then looking at 1991 to 1992 P(1) = 5.74 P(2) = 6.17 so to check 5.74 x 1.08 = 6.199 so it appears the growth rate is approx 8%...

OpenStudy (campbell_st):

so for me... I'd probably check A to C. and D the population won't decrease... just gets to a max value... so I'd leave that out... after looking at it again

OpenStudy (campbell_st):

hope it all helps

OpenStudy (anonymous):

thank you!:) @campbell_st

OpenStudy (campbell_st):

glad to help.... and I hope it was correct...